QUANDALEDINGL3:

Solve for x and show your steps. Is the solution extraneous? Check your work to show how you determined if the solution is extraneous or not. sqrt 2x minus 7 equal 1

jhonyy9:

\[\sqrt{2x -7} = 1\] in this way ? or \[\sqrt{2x} -7 = 1 \] ???